OpenStudy (anonymous):

Name four angles whose tangent equals 0.

OpenStudy (anonymous):

Look at the graph of a tan function https://www.google.com/search?q=tan(x)&oq=tan(x)&aqs=chrome..69i57j0l5.1589j0j7&sourceid=chrome&espv=210&es_sm=93&ie=UTF-8 when does it seem like it goes to 0?

OpenStudy (anonymous):

Hint: it's periodic, meaning that once you find the distance between two points of 0, that number multiplied by a integer (n) will give you another 0

OpenStudy (anonymous):

It's 0 at (0,0), but look to the left or right to the next point where the graph is 0, how far is it from 0?

OpenStudy (anonymous):

45°, 135°, 405°, 495°

OpenStudy (anonymous):

0°, 180°, 360°, 540°

OpenStudy (anonymous):

Did you take a look at the graph of tan(x)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

90°, 450°, 810°, 1170°?

OpenStudy (anonymous):

Ok, now how far is the next 0 from either side?

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

So now you know that pi(n) where n is any integer gives you an integer. And I know you know how to convert pi radians to degrees so that'll be whatever it is in degrees(n) and look at your answers accordingly

OpenStudy (anonymous):

so 3.14 divided by the answer

OpenStudy (anonymous):

naw i"ll show you

OpenStudy (anonymous):

To convert from radians to degrees or vice versa Rad-->Deg \[\Huge Rad~value~\times\frac{180}{\pi}\] Deg-->Rad\[\Huge Deg~value~\times\frac{\pi}{180}\]

OpenStudy (anonymous):

With that being said, we have pi radians (the distance) an =d if we convert to deg\[\Huge \pi~\times\frac{180}{\pi}=180\]

OpenStudy (anonymous):

so 180n are all the values for when tan(x) is 0 n=any integer (......-2,-1,0,1,2,.....)

OpenStudy (anonymous):

0°, 180°, 360°, 540°

OpenStudy (anonymous):

@raijuenea1 did you get all that?

OpenStudy (anonymous):

yes i got it thank you